write the equasion of the line which passes through point (2,1) and is parallel to the line with equasion y=5/2x+6
Answer by checkley75(3666)

A PARALLEL LINE NEEDS TO HAVE THE SAME SLOPE. THUS THE SLOPE IS 5/2. NOW SUBSTITUTE (2,1) FOR X & Y IN THE QUADRATIC EQUATION Y=mX+b
1=5/2*2+b
1=5+b
b=1-5
b=-4 Y INTERCEPT.
THEREFORE THE EQUATION OF THIS PARALLEL LINE PASSING THRU (2,1) IS
Y=5/2X-4